Circular No. 6627

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997bs IN NGC 3627 R. R. Treffers, C. Y. Peng, and A. V. Filippenko, University of Cal-

ifornia at Berkeley; and M. W. Richmond, Princeton University, report the discovery on Apr. 15 UT of a supernova (R = 17.0 ± 0.2) located ∼ 13′′ west and 67′′ south of the nucleus of NGC 3627 (α = 11^{h}20^{m}15^{s}.1, δ = +12^{o}59′22′′, equinox 2000.0). This is the first supernova found as part of the Lick Observatory Supernova Search (the direct successor of the former search at Leuschner Observatory), which is conducted with the fully automated 0.8-m Katzman Automatic Imaging Telescope. An Apr. 10 frame (limiting mag 17.6) shows no image at the location of SN 1997bs. A CCD spectrogram (range 320–800 nm, resolution 0.7 nm), obtained with the Lick 3-m Shane reflector on Apr. 16 by Filippenko, A. J. Barth, and A. M. Gilbert (also of Berkeley), shows that the object is a peculiar type-II supernova (formally known as type IIn), dominated by relatively narrow (FWHM = 1000 km/s) Balmer emission lines on a featureless continuum. There are also many weaker Fe II emission lines. SNe 1973R and SN 1989B also occurred in NGC 3627.

GRS 1737–31 Y. Ueda, T. Dotani, S. Uno, H. Inoue, and F. Nagase, Institute of

Space and Astronautical Science; Y. Maeda and M. Sakano, University of Kyoto; P. Durouchoux, Centre d’Etudes de Saclay; and J. H. Swank, Goddard Space Flight Center, NASA, report on behalf of the ASCA team: “The hard-x-ray transient GRS 1737–31 (IAUC 6599) was observed with ASCA during Apr. 2.37–2.94 UT (net exposure 27 000 s). The source was detected with an average intensity of 6 mCrab (2–10 keV) at α = 17^{h}40^{m}09^{s}, δ = −31^{o}02′.4 (equinox 2000.0; error radius 0′.5), which is consistent with the BeppoSAX position determination (IAUC 6606). The energy spectrum is well represented by a power law with photon index 1.6 ± 0.1, with a hydrogen column density of (6 ± 1) × 10^{22} cm^{−2}. The large absorption is consistent with the location close to the galactic center. Assuming the source distance to be 8500 pc, the 2–10-keV x-ray luminosity was 1.6 × 10^{36} erg/s after correcting for absorption.”

SUPERNOVA 1997br IN ESO 576-40 Corrigendum. On IAUC 6623, line 14, for Ca II, Mg II, Si II, and S II

read Si II and Fe

1997 April 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green